OpenStudy (anonymous):

Can you show me how to do this? I don't understand. '10= c/3 - 4 + c/6'

OpenStudy (jdoe0001):

do you know what the LCD is?

OpenStudy (anonymous):

Not really, no.

OpenStudy (anonymous):

@jdoe0001

OpenStudy (anonymous):

Oh! Yeah! I do.

OpenStudy (jdoe0001):

ok.... so....lemme do it by .... getting rid of the denominator.... what would be the result of \(\bf 3\times \cfrac{2}{3} \)

OpenStudy (jdoe0001):

ohhh ok, let's use the LCD then.... what's the LCD of the numbers 3 and 6 ?

OpenStudy (anonymous):

6

OpenStudy (jdoe0001):

okk...let us use that then \(\bf 10 = \cfrac{c}{3}-4+\cfrac{c}{6}\implies 10+4=\cfrac{c}{3}+\cfrac{c}{6}\qquad \color{blue}{LCD =6\quad thus}\\ \quad \\ 14=\cfrac{2c+c}{6}\implies 14=\cfrac{3c}{6}\quad \textit{multiply by 6 both sides}\\ \quad \\ 6\times 14=\cancel{6}\times \cfrac{3c}{\cancel{6}}\implies 84=3c\implies \cfrac{84}{3}=c\)

OpenStudy (anonymous):

oh; I just figured this all out with my mom, sorry! I did that before I read this...

OpenStudy (anonymous):

@jdoe0001

OpenStudy (jdoe0001):

np